Silver Company makes a product that is very popular as a Mother’s Day gift. Thus, peak sales occur in May of each year, as shown in the company’s sales budget for the second quarter given below:

April May June Total
Budgeted sales (all on account) $450,000 $650,000 $250,000 $1,350,000

From past experience, the company has learned that 25% of a month’s sales are collected in the month of sale, another 60% are collected in the month following sale, and the remaining 15% are collected in the second month following sale. Bad debts are negligible and can be ignored. February sales totaled $380,000, and March sales totaled $410,000.

Required:

1. Prepare a schedule of expected cash collections from sales, by month and in total, for the second quarter.

SCHEDULE OF EXPECTED CASH COLLECTIONS
APRIL MAY JUNE TOTAL
FEB. SALES
MARCH SALES
APRIL SALES
MAY SALES
JUNE SALES
TOTAL CASH COLLECTIONS

2. What is the accounts receivable balance on June 30th?
Answer #1

Answer

  • Requirement 1

Schedule of Expected Cash Collections

April

May

June

Total

February sales

$57,000

$57,000

March sales

$246,000

$61,500

$307,500

April sales

$112,500

$270,000

$67,500

$450,000

May sales

$162,500

$390,000

$552,500

June sales

$62,500

$62,500

Total cash collections

$415,500

$494,000

$520,000

$1,429,500

--Working

April

May

June

February sales

=380000*15%

March sales

=410000*60%

=410000*15%

April sales

=450000*25%

=450000*60%

=450000*15%

May sales

=650000*25%

=650000*60%

June sales

=250000*25%

  • Requirement 2

Account receivables on June 30:

From May Sale (650000 x 15%)

$97,500

From June sale (250000 x 75%)

$187,500

Total = ANSWER

$285,000